# Programmable Cognitive Bias in Social Agents

### v2

@@ 5.2 Testbed Extensibility
Specifically, open-ended means that the testbed can be continuously expanded in three main ways: (1) by incorporating new types of cognitive biases based on classic social science experiments—such as the Ben Franklin Effect (Franklin, 2005), the Halo Effect (Nisbett & Wilson, 1977), and others; (2) by introducing additional experimental paradigms for each bias type, for example, extending the Bandwagon Effect to include look-up studies—where individuals are influenced to look up simply because a group is doing so (Milgram et al., 1969)—to capture new aspects of social influence; and (3) by enriching the scene-adjustable placeholders within each experimental scenario to generate a wider range of prompt variations. (4) by expanding the methods for quantifying agent responses. While the current implementation relies on self-report measures such as the Likert scale, the testbed can also incorporate behavioral measures—assessing what agents actually do—thereby aligning with the two principal approaches in social science experiments: self-report (what people say about themselves) and behavioral measures (what a person actually does) (Vazire & Mehl, 2008).
@@
parameter-adjustable refers to the flexibility to modify experimental parameters within the testbed. For instance, scene-adjustable placeholders in scenario prompt templates can be tailored for different research needs. Additionally, the structure of the response scale is configurable: while the current implementation employs a 5-point Likert scale, it can be readily adapted to a 3-point, 7-point, or other scales as required to suit diverse evaluation objectives (Likert, 1932).
6 Behavioral Regulation Engine
The Behavioral Regulation Engine operationalizes the systematic behavioral control by targeting desired Cognitive Bias Index (CBI) levels. We provide comprehensive control coverage by intervening across the three fundamental spaces of a large language model: the input space ($x$), the activation space ($h$), and the parameter space ($\theta$). Formally, we model LLM generation as a function:
$y=f_{\theta}(h(x))$ (5)
where $x$ represents the input, $h(x)$ represents the hidden states/activations derived from input $x$, $\theta$ represents the model parameters, and $y$ is the generated output.
Each of the three alignment methods presents distinct trade-offs. In the input space, prompt engineering is a training-free method for guiding LLMs (Rath et al., 2025; Zhuo et al., 2024), but its effectiveness can be limited by the pitfalls of prompt design (Zamfirescu-Pereira et al., 2023) and the extreme sensitivity to minor prompt variations (Sclar et al., 2024). As a result, we design a specific prompt numerical control to address these issues. In the activation space, representation engineering manipulates internal activations during inference (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024) to control personality (Chen et al., 2025), preferences (Li et al., 2025), harmlessness (Arditi et al., 2024), and alignment (Shi et al., 2024). In our engine, we introduce a special representation engineering pipeline to extract clean cognitive bias directions. In the parameter space, fine-tuning methods like RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) are commonly used yet require substantial training data, and we introduce a bias finetuning pipeline for lightweight behavioral control utilizing LoRA (Hu et al., 2022) with task vectors (Ilharco et al., 2023). (See Appendix A.2 for discussion on control methods.) Below, we present the specific design for each space:
@@ 6 Behavioral Regulation Engine
Figure 6: Behavioral Regulation Engine. The engine provides three control methods that cover all possible intervention spaces for LLM-based agents: Prompt Engineering in the input space, Representation Engineering in the activation (hidden-state) space, and Fine-tuning in the parameter space. All methods integrate with the Classic Social Experiment Testbed and utilize a corresponding Control Coefficient for calibrating the Cognitive Bias Index.
6.1 Input Space: Prompt Numerical Control
@@
A key challenge in input space control lies in extracting a clear bias signal due to the ambiguity of qualitative adverbs in instructions (e.g., “very,” “moderately”), which LLMs interpret inconsistently (Chen et al., 2025). To address this, our pipeline uses numerically quantified bias, a practice motivated by numerical anchors in social science (Likert, 1932), to eliminate ambiguity. Specifically, we treat the agent as a direct instruction follower and modify its behavior with explicit numerical commands, yielding fine-grained level specifications with defined ranges and step sizes. Below is the prompt template for input-space control:
@@
(where 0%=[No Codified Behavior Patterns], 100%=[Maximum Codified
Behavior Patterns]). Each level increases bias by exactly 5%. Act
with this precise level [L]% of [Cognitive Bias Type].
@@
6.2 Activation Space: Representation Engineering Control
@@
Our activation space control regulates the agent’s behavior in a context-aware manner during inference (Li et al., 2023b). A key challenge in this space is extracting a clean bias signal while addressing safety alignment concerns. Modern RLHF-trained LLMs (Ouyang et al., 2022) undergo safety alignment finetuning and thus often interpret explicit bias directives (e.g., “you always trust authority figures”) as unsafe, responding with refusals (Arditi et al., 2024). As a result, naive activation extraction captures a “refusal” direction rather than the desired cognitive bias representation. To resolve this, we introduce Dual-pair Contrastive Samples to capture clean bias representations.
@@
We first generate diverse scenarios containing bias-eliciting contexts using state-of-the-art models (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). By considering only the most extreme responses (e.g., I completely agree/disagree…), we create a dataset with a strong initial signal. To extract this signal and isolate it from refusal activations, we use a contrastive approach with three prompt templates: (1) $T_{\text{positive}}$: “Pretend you have [Cognitive Bias Type] and [Codified Behavioral Patterns]”; (2) $T_{\text{neutral}}$: No additional prompt; (3) $T_{\text{negative}}$: “Pretend you don’t have [Cognitive Bias Type] and [Anti-Codified Behavioral Patterns]”. For a scenario $s_{i}$ and response $r$, we construct two contrastive pairs:
$\text{Pair}_{1}$ $:(T_{\text{positive}},s_{i},r)\text{ vs. }(T_{\text{neutral}},s_{i},r)$ (6)
$\text{Pair}_{2}$ $:(T_{\text{neutral}},s_{i},r)\text{ vs. }(T_{\text{negative}},s_{i},r)$ (7)
For each pair and token position $t$, we compute a difference vector $\Delta h_{i,t}=h^{+}_{i,t}-h^{-}_{i,t}$. After applying Principal Component Analysis (PCA) on all difference vectors, the first principal component of these difference vectors, $v_{\text{bias}}$, serves as our extracted clean bias signal, while the refusal vectors will be in the rest of the orthogonal directions. $v_{\text{bias}}$ is then used to control the agent at inference time using a Control Coefficient $\lambda$. Below, we provide two representation engineering variations:
@@
$h^{\prime}_{t}=h_{t}+\lambda\cdot v_{\text{bias}}$ (8)
RepE Projection Control: This adaptive method makes the intervention context-aware by scaling its strength based on the alignment between the hidden state and the bias vector. Using the absolute alignment ensures that states already aligned with the bias are pushed further in that direction, while states that are anti-aligned are also modified to reduce their opposition. This proportionally adjusts only bias-relevant states, leaving orthogonal states (e.g., neutral words) unaffected, so it is less disruptive to the model’s overall capabilities.
$h^{\prime}_{t}=h_{t}+\lambda\cdot|\langle h_{t},v_{\text{bias}}\rangle|\cdot v_{\text{bias}}$ (9)
6.3 Parameter Space: Fine-Tuning Control
@@
Parameter Space control involves permanently editing the model’s understanding of a specific bias. In order to make it lightweight, our process begins by using the curated bias-rich scenarios to create training data (the same data used in activation space control) for two separate LoRA models (Hu et al., 2022). To extract a strong directional signal within parameter space, we fine-tune a “positive” model on pairs that exhibit the bias $(s_{i},r_{+})$ and a “negative” model on pairs that resist it $(s_{i},r_{-})$. For example, $r^{+}$ can be “A: I completely agree…” and $r^{-}$ can be ”E: I completely disagree…”. The bias signal is then extracted by computing task vectors (Ilharco et al., 2023) from the resulting parameters:
$v_{\text{task}}$ $=\theta_{\text{positive}}-\theta_{\text{negative}}$ (10)
@@
This contrastive approach isolates a pure directional vector representing the behavioral tendency towards the bias. This signal is then used to control the agent by modifying the original model weights with a Control Coefficient $\lambda$:
$\theta_{\text{controlled}}=\theta_{\text{original}}+\lambda\cdot v_{\text{task}}$ (11)
7 Technical benchmark
7.1 Experimental Setup
Models. We evaluate two categories of large language models. The first category includes open-source models with full internal access: Llama-3.1-8B-Instruct (Llama), Mistral-7B-Instruct-v0.3 (Mistral), Qwen3-8B (Qwen), and DeepSeek-R1-0528-Qwen3-8B (DeepSeek). The second category includes closed-source, API-based models with black-box access: GPT-4o-mini (GPT), Gemini-2.5-Flash (Gemini), Claude-Sonnet-4 (Claude), and Mistral Nemo (Mistral-n). We use a sampling temperature of $T=0.7$, a standard value chosen to balance plausible response diversity with coherence. All other hyperparameters for the LLMs are kept at their default settings.
Evaluation Metrics. We assess control quality across the three properties established in Section 4: monotonicity, smoothness, and expressiveness. Let $x_{1},x_{2},\ldots,x_{n}$} denote the control coefficient settings for a method and $y_{1},y_{2},\ldots,y_{n}$} the corresponding Cognitive Bias Index (CBI) values. We formulate the metrics as follows:
  • Monotonicity: evaluates whether stronger interventions consistently move outputs in the intended direction. Soft monotonicity allows minor deviations by comparing actual sequences to ideal monotonic ones, which is measured by NDCG: $\text{NDCG}=\frac{\text{DCG}}{\text{IDCG}},\ \text{DCG}=\sum_{i=1}^{n}\frac{rel_{i}}{\log_{2}(i+2)}$, with $rel_{i}=\exp(\alpha\cdot\frac{y_{actual}[i]-y_{ideal}[i]}{y_{range}})$. Strict monotonicity (Spearman’s $\rho$) measures exact rank preservation: $\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}$, where $d_{i}$ is the rank difference between $x_{i}$ and $y_{i}$.
  • Smoothness: quantifies the stability of control adjustments. First-order smoothness ($\Delta^{1}$) captures local stability as the mean absolute difference of adjacent bias values: $\Delta^{1}=\text{Mean}(|y_{i+1}-y_{i}|)$. Second-order smoothness ($\Delta^{2}$) captures trend continuity as the mean absolute difference of consecutive first-order changes: $\Delta^{2}=\text{Mean}(|(y_{i+1}-y_{i})-(y_{i}-y_{i-1})|)$.
  • Expressiveness: measures the effective control range, defined as the difference between maximum and minimum observed bias levels: $\text{Expressiveness}=\max(y_{i})-\min(y_{i})$.
7.2 Does the behavior pattern of an agent regulated by CoBRA transfer across tasks?
A central hypothesis behind CoBRA is that ”aligning an agent’s behavior with target profiles in a set of classical experiments will reliably induce the corresponding cognitive biases, which will then generalize and manifest across a broader range of tasks”. We study this transferrability problem by testing CoBRA across different classic social experimental paradigms for Framing Effect.
Setup. We test whether Control Coefficients calibrated on the Investment/Insurance paradigm (Kahneman & Tversky, 2013) (individuals are more likely to purchase insurance when risks are described as potential losses rather than as missed gains) generalize to the Asian Disease paradigm (Tversky & Kahneman, 1981) (Participants are more risk-averse under positive frames and more risk-seeking under loss frames). We select 10 diverse agent personas from the persona pools in Generative Agents (Park et al., 2023), with varied specifications. For each persona, we sweep over all possible Control Coefficients. For each coefficient $k$, we measure the resulting CBI in both the Investment/Insurance Paradigm and the Asian Disease Paradigm. This generates paired CBI sequences, $(y_{k}^{\text{invest}},y_{k}^{\text{asian}})$, for each persona across the two experiments. We also evaluate CoBRA using the Behavioral Regulation Engine, activating only one method at a time.
Figure 7: Reproducibility experiments across paradigms. The x-axis represents the Cognitive Bias Index measured in the Investment/Insurance Paradigm, while the y-axis shows the Cognitive Bias Index using the same control coefficients in the Asian Disease Paradigm. Each thin line represents a unique agent persona, and the bold line is the average of them. Control Coefficients calibrated on the Investment/Insurance Paradigm produce a linearly correlated CBI response in the Asian Disease Paradigm.
Results. The averaged trajectory curves demonstrate complete linear correlation between paradigms (Fig. 7). This linear relationship indicates that control strength transfers proportionally: agents exhibiting high framing susceptibility in Investment/Insurance Paradigm show correspondingly high susceptibility in Asian Disease Paradigm. The CBI is generalizable across experimental paradigms.
7.3 Reproducibility
Reproducible bias control across different model configurations is fundamental for scientific rigor in social science simulation research. We ensure that when specifying a particular level of cognitive bias, the LLM-based agent will produce equivalent behavioral tendencies regardless of the foundation model, sampling temperature, or whether reasoning is used. Without this reproducibility guarantee, results can’t be reliably replicated, compared, or validated—undermining the credibility of LLM-based social research.

### v3

@@ 5.2. Testbed Extensibility
Specifically, open-ended means that the testbed can be continuously expanded in four main ways: (1) by incorporating new types of cognitive biases based on classic social science experiments—such as the Ben Franklin Effect (Franklin, 2005), the Halo Effect (Nisbett and Wilson, 1977), and others; (2) by introducing additional experimental paradigms for each bias type, for example, extending the Bandwagon Effect to include look-up studies—where individuals are influenced to look up simply because a group is doing so (Milgram et al., 1969)—to capture new aspects of social influence; and (3) by enriching the scene-adjustable placeholders within each experimental scenario to generate a wider range of prompt variations. (4) by expanding the methods for quantifying agent responses. While the current implementation relies on self-report measures such as the Likert scale, the testbed can also incorporate behavioral measures—assessing what agents actually do—thereby aligning with the two principal approaches in social science experiments: self-report (what people say about themselves) and behavioral measures (what a person actually does) (Vazire and Mehl, 2008).
@@
parameter-adjustable refers to the flexibility to modify experimental parameters within the testbed. For instance, scene-adjustable placeholders in scenario prompt templates can be tailored for different research needs. Additionally, the structure of the response format is configurable: While the current implementation uses a 5-point multiple-choice format, it can be easily adapted to 3-point, 7-point, or other scales to suit different evaluation needs.
@@
A system diagram of the Behavioral Regulation Engine, illustrating three complementary intervention mechanisms for controlling LLM-based agent behavior: prompt engineering in the input space, representation engineering in the activation or hidden-state space, and parameter-level control via fine-tuning. Each control method interfaces with the Classic Social Experiment Testbed and uses a control coefficient to calibrate the agent’s Cognitive Bias Index.
Figure 5. Behavioral Regulation Engine. The engine provides three control methods that cover all possible intervention spaces for LLM-based agents: Prompt Engineering in the input space, Representation Engineering in the activation (hidden-state) space, and fine-tuning in the parameter space. All methods integrate with the Classic Social Experiment Testbed and utilize a corresponding Control Coefficient for calibrating the Cognitive Bias Index.
6. Behavioral Regulation Engine
The Behavioral Regulation Engine operationalizes the systematic behavioral control. To meet diverse usage needs, we intervene across all three fundamental spaces of a large language model (input space, activation space, and parameter space), ensuring flexibility in behavioral control.
Each of these spaces requires distinct alignment methods with specific trade-offs. In the input space, prompt engineering is a training-free method for guiding LLMs (Rath et al., 2025; Zhuo et al., 2024), but its effectiveness can be limited by the pitfalls of prompt design (Zamfirescu-Pereira et al., 2023) and the extreme sensitivity to minor prompt variations (Sclar et al., 2024), with risks of instruction-following artifacts that reduce control precision and smoothness. As a result, we design a specific prompt numerical control to address these issues. In the activation space, representation engineering manipulates internal activations during inference (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024) to control personality (Chen et al., 2025), preferences (Li et al., 2025), harmlessness (Arditi et al., 2024), and alignment (Shi et al., 2024), though extraction quality depends on isolating clean signals from confounding factors like safety refusals. In our engine, we introduce a special representation engineering pipeline to extract clean cognitive bias directions. In the parameter space, fine-tuning methods like RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) are commonly used yet require substantial training data, with potential for overfitting to training scenario patterns. We introduce a bias fine-tuning pipeline for lightweight behavioral control utilizing LoRA (Hu et al., 2022) with task vectors (Ilharco et al., 2023), with specific design choices to mitigate these concerns. Below, we present the specific design for each space:
6.1. Input Space: Prompt Numerical Control
@@
A key challenge in input space control lies in extracting a clear bias signal due to the ambiguity of qualitative adverbs in instructions (e.g., “very,” “moderately”), which LLMs interpret inconsistently (Chen et al., 2025). To address this, our pipeline uses numerically quantified bias, a practice motivated by numerical anchors in social science (Likert, 1932), to mitigate ambiguity. Specifically, we treat the agent as a direct instruction follower and modify its behavior with explicit numerical commands, yielding fine-grained level specifications with defined ranges and step sizes (See validation in Appendix C.1). Prompt template for input-space control:
@@
(where 0%=[No Codified Behavior Patterns],
100%=[Maximum Codified Behavior Patterns]).
Each level increases bias by exactly 5%. Act with this precise
level [L]% of [Cognitive Bias Type].
@@
6.2. Activation Space: Representation Engineering Control
@@
Our activation space control regulates the agent’s behavior in a context-aware manner during inference (Li et al., 2023b). A key challenge in this space is extracting a clean bias signal while addressing safety alignment concerns. Modern RLHF-trained LLMs (Ouyang et al., 2022) undergo safety alignment fine-tuning and thus often interpret explicit bias directives (e.g., “you always trust authority figures”) as unsafe, responding with refusals (Arditi et al., 2024). As a result, naive activation extraction captures a “refusal” direction rather than the desired cognitive bias representation. To resolve this, we introduce Dual-pair Contrastive Samples to isolate clean bias representations while mitigating confounds such as refusals.
Extracting Bias Representations. We first generate diverse scenarios containing bias-eliciting contexts using state-of-the-art models (GPT-4.1, Claude-Sonnet-4, Gemini-2.5-Pro, and Grok-4). By considering only the most extreme responses (e.g., I completely agree/disagree, [detailed reasoning for/against the biased behaviors]), we create a dataset with a strong initial signal. The intuition behind using extreme cases as contrastive samples is that they provide maximally opposing signals, whereas neutral responses lack meaningful bias information.
@@
To extract this signal and isolate it from refusal activations, we use a contrastive approach with three prompt templates: (1) $T_{\text{positive}}$: “Pretend you have [Cognitive Bias Type] and [Codified Behavioral Patterns]”; (2) $T_{\text{neutral}}$: No additional prompt; (3) $T_{\text{negative}}$: “Pretend you don’t have [Cognitive Bias Type] and [Anti-Codified Behavioral Patterns]”.
For each scenario $s_{i}$ and response, we compute hidden state representations $h_{i,t}$ extracted from the model’s intermediate layers at token position $t$. Using the three contrastive prompts described earlier ($T_{\text{positive}},T_{\text{neutral}},T_{\text{negative}}$), we obtain the hidden states $h_{i,t}^{+},h_{i,t}^{0},$ and $h_{i,t}^{-}$, corresponding to the positive, neutral, and negative conditions for each token $t$. We then construct two difference vectors for each token position $t$:
(1) $\Delta h_{i,t}^{(1)}$ $=h_{i,t}^{+}-h_{i,t}^{0},$
(2) $\Delta h_{i,t}^{(2)}$ $=h_{i,t}^{0}-h_{i,t}^{-}.$
These difference vectors encode the changes in activation caused by the presence or absence of the bias signal. For example, $\Delta h_{i,t}^{(1)}$ captures the shift in activation when moving from a neutral context to one where the bias is explicitly encouraged, while $\Delta h_{i,t}^{(2)}$ captures the shift when discouraging the bias instead.
The intuition here is that these differences isolate the bias-relevant signal by subtracting out shared components of unrelated activations (e.g., shared context). To further distill the dominant bias direction, we apply Principal Component Analysis (PCA) on all difference vectors $\Delta h_{i,t}$ to get rid of confounding factors. The first principal component, $v_{\text{bias}}$, represents the most consistent and robust bias signal over all token positions and contrastive pairs.
Applying Inference-Time Control. Having extracted the clean bias signal $v_{\text{bias}}$, we then decided which layers to control. Rather than pre-selecting fixed layers, we adopted an empirical, data-driven approach. We performed a layer-wise evaluation on a test set of our contrastive samples. For each layer, we used the sign of the inner product of $h_{i,-1}$(i.e., hidden states of last token) and $v_{bias}$, as a linear classifier to predict whether a given context was bias-eliciting (+1) or bias-resisting (-1). The layers with the highest classification accuracy were selected. Control interventions were then applied to the top 15 layers exhibiting the highest predictive accuracy, as these layers contain the most robust and influential representations of the target bias. (See Appendix C for further justification)
We use $v_{bias}$ on selected layers to control the agent’s behavior during inference via a Control Coefficient $\lambda$, which modulates its influence on the model’s activations. Linear Control provides a simple, straightforward approach for uniform bias injection, while Projection Control enables context-aware adjustments for complex scenarios. Here, $h_{t}$ denotes the hidden state during inference, adjusted using $v_{\text{bias}}$ to guide the model’s behavior:
@@
(3) $h^{\prime}_{t}=h_{t}+\lambda\cdot v_{\text{bias}}$
The motivation for Linear Control lies in its simplicity and effectiveness. By directly adding the bias vector $v_{\text{bias}}$, it uniformly injects the bias signal across all activations, making it highly effective for achieving a large control range. However, its context-independent nature may over-correct neutral contexts, reducing the smoothness of control.
RepE Projection Control: This adaptive method makes the intervention context-aware by scaling its strength based on the alignment between the hidden state and the bias vector:
(4) $h^{\prime}_{t}=h_{t}+\lambda\cdot|\langle h_{t},v_{\text{bias}}\rangle|\cdot v_{\text{bias}}$
Projection Control is introduced to address the limitations of Linear Control by introducing context-awareness. The intervention is dynamically scaled by the magnitude of the alignment between the hidden state and the bias vector, ensuring that adjustments are applied proportionally to how relevant the current context is to the bias (i.e., for an unrelated context, the intervention is minimal). This context-aware approach results in better smoothness, though it offers a more constrained control range.
Ensuring Genuine Behavioral Change. Activation-space methods, by their design, do not have overfitting issues. Our contrastive extraction method uses subtraction on hidden states to cancel shared features between contrastive pairs, such as question and answer formats, thereby isolating the underlying bias direction. In addition, we modify context-dependent representations rather than surface embeddings, which changes the model’s internal computational mechanisms and ultimately leads to behavioral changes (Conmy et al., 2023; Templeton et al., 2024).
6.3. Parameter Space: Fine-tuning Control
@@
Extracting Bias Task. Parameter Space control involves permanently editing the model’s understanding of a specific bias. In order to make it lightweight, our process begins by using the curated bias-rich scenarios to create training data (the same data used in activation space control) for two separate LoRA models (Hu et al., 2022). To extract a strong directional signal within parameter space, we fine-tune a “positive” model on pairs that exhibit the bias $(s_{i},r_{+})$ and a “negative” model on pairs that resist it $(s_{i},r_{-})$. For example, $r^{+}$ can be “A: I completely agree, [detailed reasoning for the biased behaviors]” and $r^{-}$ can be ”E: I completely disagree, [detailed reasoning against the biased behaviors]”. The bias signal is then extracted by computing task vectors (Ilharco et al., 2023) from the resulting parameters:
(5) $v_{\text{task}}$ $=\theta_{\text{positive}}-\theta_{\text{negative}}$
Applying Weight-Based Control
@@
This contrastive approach isolates a pure directional vector representing the behavioral tendency towards the bias (not only on choices). This signal is then used to control the agent by modifying the original model weights with a Control Coefficient $\lambda$:
(6) $\theta_{\text{controlled}}=\theta_{\text{original}}+\lambda\cdot v_{\text{task}}$
Ensuring Genuine Behavioral Change. Fine-tuning methods have the risk of overfitting on specific test formats. We reduce overfitting by training on diverse, out-of-domain data generated by multiple models. Additionally, the training data contains reasoning-based responses that extend beyond simple answer choices, ensuring the learned bias direction generalizes beyond format-specific patterns.
7. Evaluation
CoBRA aims to (1) improve the reproducibility of agent behaviors across models and (2) enable explicit, quantitative control over those behaviors. This section reports experiments evaluating both aims.
7.1. Reproducibility
As shown in Section 3, the same specification yields inconsistent behaviors across different models. As a result, others may be unable to reproduce the simulation because access to specific foundation models can change over time. In this experiment, we evaluate whether CoBRA improves behavioral consistency across models, across sampling temperatures, and with or without reasoning enabled.
7.1.1. Across Models
We evaluated cross-model behavioral consistency using 4 popular open-source foundation models (Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, DeepSeek-R1-0528-Qwen3-8B, Qwen3-8B). Given a foundation model, we first used CoBRA to create agents with varying levels of specified bias. Our testbed (Table 1) contains four cognitive biases, each with two built-in experiment paradigms. For each bias, we used one experiment paradigm to calibrate the agent’s behavior and measured the agent’s responses in the other paradigm.
We then characterized behavioral consistency by measuring the variance in agents’ responses across different models in the same experiment. Each experiment paradigm contains between 15 and 75 questions, depending on the number of factors involved. For each question, we collected the model’s internal probability distribution over response options, directly extracted from the output logits via softmax. We then computed the variance by averaging the absolute differences between pairwise responses from different models.
Baselines. We included two baselines to contextualize the effect of CoBRA. The first, No control, captures response variance across all models without any prompt intervention. The second, Control, measures response variance when applying an aggressive prompt-based bias injection (i.e., “Pretend you have completely [bias type]”).
Results. Figure 6 presents the response variable across four types of cognitive bias. All CoBRA methods significantly reduce cross-model variance relative to both baselines (all $p<0.001$; mean Cohen’s d=2.41d=-2.41 vs. No Control, d=9.20d=-9.20 vs. Control). RepE Projection yields the lowest mid-range variance with minimal deviation across the coefficient spectrum (coefficient of variation $4\%$).
Refer to caption A line plot comparing cross-model variance under different control conditions. The x-axis represents the specified bias level in a calibrating paradigm, and the y-axis shows the variance of agent behavior across models in a target paradigm. The curves indicate that CoBRA substantially reduces cross-model variance compared to baseline and no-control conditions, with shaded bands representing variability.
Figure 6. CoBRA methods significantly reduce variance compared to both baselines (all $p<0.001$, mean Cohen’s d=2.41d=-2.41 vs No Control, d=9.20d=-9.20 vs Control). The x-axis represents the specified amount of bias in the calibrating paradigm, while the y-axis shows the variance across models in the target paradigm. Shaded regions indicate $\pm 1$ SE.
7.1.2. Across Sampling Temperatures
We evaluated cross-temperature behavioral consistency by measuring agent behavior as we varied the temperature over $T\in\{0.1,0.2,\dots,1.0\}$. At each temperature, we sweep the control coefficients and then observe whether the responses to the experiment paradigm are stable. In earlier experiments, we observed no significant differences among models or experimental groups. So we used the Mistral as the representative model and the Milgram paradigm to test the agents we created.
Results. Figure 7 illustrates how the agent’s responses change when we adjust the control coefficient (x-axis) and temperature (different lines). Agents regulated by CoBRA are mostly insensitive to temperature changes. We performed pairwise equivalence tests (Two One-Sided Tests) (Schuirmann, 1987) across all $\binom{10}{2}=45$ temperature pairs, with a threshold of 0.5 (12.5% of the 1-5 Likert scale).
The equivalence testing proves all temperature curves are equivalent (45/45 pairs, all $p<0.05$), with temperature contributing only 0.5-1.2% of total behavioral variance. Representation engineering methods show an even smaller gap across temperatures. The results indicate that the Control Coefficient we obtained under one temperature can be directly applied in experiments using a different temperature, demonstrating the reproducibility across sampling temperatures.
Refer to caption A line plot examining the effect of sampling temperature on bias control. The figure shows Cognitive Bias Index values as a function of the control coefficient for multiple temperature settings, ranging from low to high randomness. The tightly clustered curves with overlapping variability bands indicate that changes in temperature do not substantially affect the bias control trajectory, demonstrating reproducible behavior across sampling temperatures.
Figure 7. Reproducibility experiments across sampling temperatures. The x-axis is the Control Coefficient. The y-axis is the response to the target bias paradigm measured in CBI. Each colored line represents a different temperature setting, ranging from a deterministic T=0.1 to a more random T=1.0. All 45 pairwise equivalence tests prove temperature curves are equivalent ($p<0.05$). The fact that all the lines are tightly clustered and follow the same path demonstrates that the control is not affected by changes in temperature, ensuring that CoBRA ensures reproducibility results across different sampling temperatures.
7.1.3. Across Reasoning Modes
We evaluated whether the behavior of an agent regulated by CoBRA is stable when agents answer with or without explicit reasoning. Similar to Section 7.1.2, we used Mistral as the representative model and the two authority bias paradigms as the testbed, comparing Direct mode (the agent answers immediately after reading the question) and Reasoning mode (the agent first provides a brief rationale, then answers). The reasoning length is capped at 128 tokens to standardize the reasoning budget while allowing sufficient space for rationales (approximately 2-3 sentences). We repeated each question 8 times.
Results. Figure 8 shows that both methods preserve a linear correlation across paradigms, and for RepE methods, the Direct and Reasoning modes exhibit almost identical behavior. Within a single paradigm (Fig. 9), RepE Linear and RepE Projection produce nearly overlapping curves across the two modes, whereas the Prompt Numerical method is more sensitive to reasoning-mode variations. These results indicate that Control Coefficients calibrated in Direct mode transfer seamlessly to reasoning settings, and that RepE methods provide stable, continuous bias control even when the model generates intermediate reasoning. Statistical equivalence testing confirms that RepE methods maintain nearly identical behavior across both reasoning modes ($p<0.001$ for all comparisons; see Appendix D.5 for details).
Refer to caption A line plot comparing cross-paradigm reproducibility of direct and reasoning modes. The figure plots the Cognitive Bias Index measured in the Milgram Obedience paradigm on the x-axis against the index measured in the Stanford Prison paradigm on the y-axis using the same control coefficients. Separate curves represent direct and reasoning modes, with shaded bands indicating variability, showing that control coefficients generalize across paradigms for both modes.
Figure 8. Reproducibility of different reasoning modes across paradigms. The x-axis represents the Cognitive Bias Index (CBI) measured in the Milgram Obedience Paradigm, while the y-axis shows the CBI using the same Control Coefficients in the Stanford Prison Paradigm with Direct mode and Reasoning mode. The Control Coefficients generalize across different paradigms for both modes. Shaded regions represent standard $\pm 1$ SE error bands.
Refer to caption A set of control curves showing the effect of different reasoning modes within the Milgram paradigm. The plot relates the control coefficient on the x-axis to the Cognitive Bias Index on the y-axis, with separate curves for different reasoning modes. The nearly overlapping curves and shaded variability bands indicate that representation-level interventions produce stable and consistent bias control regardless of reasoning mode.
Figure 9. Control curves under different reasoning modes. The x-axis is the Control Coefficient, and the y-axis is the response to the Milgram experiment measured in CBI (Milgram, 1963). Within a single paradigm, Representation Engineering methods produce nearly identical behavior across modes, indicating stable, consistent bias control during reasoning. Shaded regions represent $\pm 1$ standard error bands.
7.2. Controllability
Another key design goal for CoBRA is that small adjustments to the index (e.g., increasing a bias level) should reliably produce small, predictable shifts in agent behavior. To understand the effectiveness of the controllability through CoBRA, we evaluate four aspects:
  • Monotonicity: Does turning the bias “up” or “down” always push the model’s answers in that same direction, without reversing unexpectedly?
  • Smoothness: Do small changes in the bias level lead to small, gradual changes in the model’s answers, rather than sudden jumps?
  • Expressiveness: How large is the gap between the least-biased and most-biased behavior that the control can produce?
  • Generalization: Does a control setting that works in one experiment still produce the intended behavior in new experiments?
7.2.1. Apparatus
In addition to the four open-source models, we also evaluated four closed-source, API-based models with black-box access: GPT-4o-mini (GPT), Gemini-2.5-Flash (Gemini), Claude-Sonnet-4 (Claude), and Mistral Nemo (Mistral-n). In open-source settings, we directly obtain option-level probability distributions from model logits to compute the CBI, whereas for closed-source, API-based models where logits are unavailable, we estimate these probabilities from the empirical frequencies of 10 samples per question. We use a sampling temperature of $T=0.7$, a standard choice that balances response diversity and coherence. All other hyperparameters are default.
Control Baseline. We design a purely implicit natural language-based control baseline. For each bias type and scenario, we instruct the model using natural-language descriptions that span multiple bias levels (i.e., from ”no authority bias” to ”extreme authority bias”) and we map to normalized control coefficients from 0 to 1. We tested this baseline control across four open-source models and collected the raw probability of response distributions across five options. Figure 10 visualizes how the response distribution shifts across control coefficients. The resulting distributions are highly uneven and hard to predict.
A grid of stacked area plots illustrating baseline control using raw coefficients across multiple bias types and experimental paradigms. Columns correspond to different cognitive biases, and rows correspond to two test paradigms. Each plot shows the probability distribution over five response levels, from strongly biased to strongly reverse-biased, as the normalized control coefficient increases. The uneven and nonmonotonic patterns indicate that direct coefficient-based control produces unpredictable and inconsistent behavioral distributions.
Figure 10. Baseline control using coefficients directly leads to an uneven distribution of behaviors. Each column represents a bias type, and each row shows one of the two test paradigms. The stacked area plots show the probability of the five response levels (from strongly biased (blue) to strongly reverse-biased (green)) as a function of the normalized control coefficient (x-axis). The resulting distributions are highly uneven and hard to predict. For example, in the Framing paradigms (third column), the bias decrease then increase—cannot produce consistent control.